OpenStudy (anonymous):

the domain of the given function. f(x) = square root of x +3 / ( x+8 ) ( x -2 )

OpenStudy (anonymous):

De swam 5 more than 3x her father last week in total = 23 -5 =18 divided by 3 =6.

OpenStudy (anonymous):

A. 3(m + 5) = 23 B. 3m + 5 = 23 C. 3m - 5 = 23 D. 5m + 3 = 23

OpenStudy (anonymous):

what does the m stand for?

OpenStudy (anonymous):

miles?

OpenStudy (anonymous):

" De swam five miles farther" + 5 5 plus what??? " than three times the distance his father swam." than 3m

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

i think you've got it